A sample to be used for medical imaging is labeled with F-18, which has a half- life of 110 min. What percentage of the original activity in the sample remains after 300min?
1
Expert's answer
2021-07-14T05:48:52-0400

Let the final activity be x

"\ln \frac{x}{100}= -\frac{0.693}{110}*300\\
\ln \frac{x}{100}= -1.89\\
\ln x - \ln 100 = -1.89\\
\ln x =2.715\\
x= 15.10 \%"